Hello,
I need help with this probelm:
Find タ-1(x)
f(x)=cube root of x-7
タ-1 (x)= ?
Thank you!

Answer by htmentor(1343)

To find the inverse, solve for x:
f(x) = y = (x-7)^(1/3)
y^3 = x - 7
x = y^3 + 7
Now switch x and y:
Then inverse is f^-1(x) = x^3 + 7